Accelerator Update for October 24, 2007 - October 25, 2007
Wednesday October 24
Operations put a Main Injector (MI) power supply (M:V105) into its DC mode (the supply and magnet stays on all the time and doesn’t ramp up and down) at 8:20 AM. This magnet appeared to have a temperature problem and experts wanted to see if this test would make it trip on a temperature indication.
At 8:48 AM, Operations reported a problem with a Pbar Debuncher TWT (D:V1TW04) that would not reset or turn on. An expert investigated, but found nothing wrong. 15 minutes later Operations reported the same problem. The expert investigated again and replaced a power supply.
At 9:49 AM, the MI power supply interlock loop tripped off due to a volts-to-ground indication at MI-20. Operators contacted an expert.
At 10:17 AM, M:V105 tripped off on an over temperature indication for the first of two times in an hour. An expert investigated.
At 11:11 AM, the MIPS Loop tripped off. Operations reset the system.
A MI expert reported at 12:22 PM that he needed to replace the power supply’s main contactor for M:V105; the magnet wasn’t causing the temperature problem. He had the power supply back in service about two hours later.
Operations put the TeV through a dry squeeze at 2:05 PM.
At 2:46 PM, the TeV ramp dumped due to a power supply failure at sector C2. Operators investigated and discovered that a technician taking a measurement had plugged into the wrong test point. Operators began recovery.
TeV experts began a study at 2:49 PM.
Operations put the TeV through a dry squeeze at 3:20 PM.
Pbar experts began a reverse proton study at 3:45 PM.
SY experts completed their study at 4:23 PM.
Recycler studies ended at 8:13 PM.
Thursday October 25
The midnight shift began TeV experts conducting studies, with Pbar conducting reverse proton studies and aperture scans, with Operations monitoring the Recycler stash, and with Sci /MiniBooNE taking beam.
Operations put the TeV through dry squeeze at 2:29 AM.
The Plans for Thursday
The current plan is for the accelerators to continue with studies and commissioning.
Linac: A RF station (LRF2) needs tuning
Main Injector: All their controls and configuration problems have been fixed. Their DCCT might need a new op amp. They want to try their aperture scan program today and conduct a coalescing study.
Antiproton Source: They have phased their momentum cooling. Their aperture scans have revealed some restrictions.
Recycler: They are working on aligning their electron beam.
Tevatron: They are continuing to work on tuning up the ramp.
Switchyard 120: They have a safety walkthrough today.
Comments and Suggestions
|last modified 10/26/2007 email Fermilab|